OpenStudy (anonymous):

what are the solutions to the nonlinear of equations below? check all that apply x^2-y^2=4 x^2/2^2+y^2/2^2=1 A.(2,2) B.(-2,-2) C.(0,2) D.(0,-2) E. (2,0) F.(-2,0)

OpenStudy (cwrw238):

what must x^2 and y^2 be equal to?

OpenStudy (anonymous):

4

OpenStudy (cwrw238):

right so what are the possible values of x and y in the list of options?

OpenStudy (anonymous):

my mistake i meant to write x^2/2^2+y^2/2^2=1

OpenStudy (cwrw238):

ahhhh

OpenStudy (cwrw238):

so we have x^2 - y^2 = 4 and x^2 y^2 --- + --- = 1 2^2 2^2 right?

OpenStudy (anonymous):

Correct

OpenStudy (cwrw238):

ok simplify the second equation by multiplying thru by 2^2 ( = 4): x^2 + y^2 = 4 ...................(1) x^x - y^2 = 4 ....................(2) now add (1) and (2) what does that give?

OpenStudy (cwrw238):

** my typo that x^2 in equation (2)

OpenStudy (anonymous):

x^2-y^2 = x^2 + y^2 +4=4 ,that's what i got

OpenStudy (anonymous):

is that right?

OpenStudy (anonymous):

I just need the answers that could either be A,b,c,d,e, or f

OpenStudy (cwrw238):

when you add the y^2's are eliminated:- y^2 + - y^2 = 0 giving 2x^2 = 8 so now its easy to find possible values of x and y

OpenStudy (anonymous):

would the final asnwer be 2,2?

OpenStudy (anonymous):

E. F.